# CollaFuse: Collaborative Diffusion Models

### v1

@@ 5 Discussion, Limitations, and Future Work
Our findings reveal that collaborative diffusion models produce images of higher quality than those generated by independently trained diffusion models using sub-datasets. In certain instances, they even surpass the performance of centrally trained diffusion models. This not only highlights the efficacy of our method but also demonstrates its capability to tackle the personalization versus generalization challenge frequently encountered in federated learning scenarios. Beyond the aspect of image fidelity, it is pivotal to consider our findings within their broader impact on information disclosure and computational efficiency in distributed machine learning frameworks. Our approach circumvents the need to share raw data or complete model updates, opting instead to share only the diffused images alongside server-side noise. Furthermore, by delegating computationally intensive tasks, our approach significantly reduces the computational load on individual clients, leveraging the potential strengths of growing foundation models on large servers. Consequently, even in cases where image fidelity decreases, clients may still favor a collaborative diffusion model for its ability to lighten their computational load, even in simple one-to-one setups. Thus, our strategy facilitates a balanced optimization of performance, privacy, and computational resources and can be tailored to individual preferences.
@@
Future Work: Our empirical evaluations are currently limited to images with $64\times 64$ resolution. An important future avenue is the collaborative training of diffusion models of larger scales, for which the computational advantage to each client further increases. Also, collaborative diffusion applications based on more general text-to-image synthesis models like Stable Diffusion are interesting to investigate. However, training such models requires high computational resources and is, therefore, out of the scope of this work. Nevertheless, future research steps include examining vulnerabilities, like backdoor attacks, in collaborative diffusion models. Identifying suitable countermeasures and analyzing their impact on image fidelity and computational efficiency is central to bringing our approach closer to its real-world application. Another intriguing avenue is the combination of our collaborative diffusion models with differentially private training algorithms to provide formal privacy guarantees for trained models [7, 9]. Similarly, it would be interesting to investigate to what extent the phenomenon of memorization in diffusion models [4, 46, 14] can also occur in collaborative approaches with separate models.
@@ 6 Conclusion
Our collaborative diffusion approach offers a novel solution to the challenges of diffusion-based generative models. By dividing the denoising process between a shared server and client models, we address performance, information disclosure, and computational concerns effectively. Our approach enables clients to outsource computationally intensive denoising steps to the server, balancing image quality without the necessity of sharing raw data. Through experiments, we demonstrate the effectiveness of collaborative training in enhancing image quality tailored to each client’s domain while reducing the number of denoising steps on the client side. These findings highlight the potential of collaborative diffusion models in advancing distributed machine learning research and development.
@@
Our source code is publicly at https://github.com/SimeonAllmendinger/collafuse to reproduce the experiments and facilitate further analysis.

### v2

@@ 5 Discussion, Limitations, and Future Work
Our findings reveal that CollaFuse produces images of higher quality than those generated by independently trained diffusion models using sub-datasets. In certain instances, they even surpass the performance of globally trained diffusion models. This not only highlights the efficacy of our method but also demonstrates its capability to tackle the personalization versus generalization challenge frequently encountered in federated learning scenarios [53]. Beyond the aspect of image fidelity, it is pivotal to consider our findings within their broader impact on information disclosure and computational efficiency in distributed machine learning frameworks. Our approach circumvents the need to share raw data or complete model updates, opting instead to share only the diffused images alongside server-side noise. Furthermore, by delegating computationally intensive tasks, our approach significantly reduces the computational load on individual clients, leveraging the potential strengths of growing foundation models on large servers. Consequently, even in cases where image fidelity decreases, clients may still favor a collaborative diffusion model for its ability to lighten their computational load, even in simple one-to-one setups. Thus, our strategy facilitates a balanced optimization of performance, information disclosure, and computational resources and can be tailored to individual preferences.
@@
Future Work: Our empirical evaluations are currently limited to four models and image resolutions. An important future avenue is the collaborative training of diffusion models of larger scales without pre-training, for which the computational advantage to each client further increases. However, training such models requires high computational resources and is, therefore, out of the scope of this work, which already utilized 11 A100 GPUs in parallel. Nevertheless, future research steps include examining vulnerabilities, like backdoor attacks, in CollaFuse. Identifying suitable countermeasures and analyzing their impact on image fidelity and computational efficiency is central to bringing our approach closer to its real-world application. Another intriguing avenue is the combination of CollaFuse with differentially private training algorithms to provide formal privacy guarantees for trained models [7, 10]. Similarly, it would be interesting to investigate to what extent the phenomenon of memorization in diffusion models [3, 49, 15] can also occur in collaborative approaches with separate models and datasets.
@@ 6 Conclusion
Our collaborative diffusion approach CollaFuse offers a novel solution to the challenges of diffusion-based generative models. By dividing the denoising process between a shared server and client models, we address performance, information disclosure, and computational concerns effectively. Our approach enables clients to outsource computationally intensive denoising steps to the server, balancing image quality without the necessity of sharing raw data. Through experiments, we demonstrate the effectiveness of collaborative training in enhancing image quality tailored to each client’s domain while reducing the number of denoising steps on the client side. These findings highlight the potential of CollaFuse in advancing distributed machine learning research and development.
@@
Our source code is publicly at https://anonymous.4open.science/r/collafuse-25ED/README.md to reproduce the experiments and facilitate further analysis.